Complete the following homework problems. Show all work (making answers for clarity sure it is legible) and circle all ProblePlease solve Q1 and Q2
Answer #1

Wu wuBc+A w2-AD Snalart = ABC+A-A+B¢0+DA Jo X Y w.wu Y=(AtD) Cac+Ā) = ABC+ Bco+ A DD U3 A C O C O C O O 0 C A A AD O O O 0 OGA Y O C O O o 1 O C C O O k-map CD AA D I | AD+ABC X AD+ABC Y - c O O Ca,cd Fdmo input inpub greater han Input inpu 2, bupud han 2 leu Qutpud mpu a bc d complmenh Truth tabk In Then X Vulue C 1 O1 a lo Y = Co+ABC+ACO AD ABDACD+BcD A B